Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 31     Entry time: 15 Feb 2023, 01:40
Author: Laurent Mercadier, Luigi Adriano 
Group: SCS 
Subject: late shift shift summary 

Achievements

  • Two color mode established with 11.5 GeV, 1.335 keV col1, 1.27 keV col2
  • Beam size characterized using Viking + online tool
  • Jet found by scanning and looking at transmitted signal on Viking

Problems

  • Issue with train Id of Chem motors --> fixed by setting proper trigger
  • Beam loss around 19:30 induced instability in the two color mode, slow degradation of intensity and stability until 23:00 when it was not possible to use the beam
    • Called PRC, decided to tune to single color, 1.3 keV for the night
  • Jet froze
ELOG V3.1.4-7c3fd00